# Logits of API-Protected LLMs Leak Proprietary Information

### 14 Mar 2024 v1

@@ Abstract
The commercialization of large language models (LLMs) has led to the common practice of high-level API-only access to proprietary models. In this work, we show that even with a conservative assumption about the model architecture, it is possible to learn a surprisingly large amount of non-public information about an API-protected LLM from a relatively small number of API queries (e.g., costing under \$1,000 for OpenAI’s gpt-3.5-turbo). Our findings are centered on one key observation: most modern LLMs suffer from a softmax bottleneck, which restricts the model outputs to a linear subspace of the full output space. We show that this lends itself to a model image or a model signature which unlocks several capabilities with affordable cost: efficiently discovering the LLM’s hidden size, obtaining full-vocabulary outputs, detecting and disambiguating different model updates, identifying the source LLM given a single full LLM output, and even estimating the output layer parameters. Our empirical investigations show the effectiveness of our methods, which allow us to estimate the embedding size of OpenAI’s gpt-3.5-turbo to be about 4,096. Lastly, we discuss ways that LLM providers can guard against these attacks, as well as how these capabilities can be viewed as a feature (rather than a bug) by allowing for greater transparency and accountability.
@@ 1 Introduction
Considering several proposals to mitigate this vulnerability, we find no obvious fix to prevent obtaining LLM images without dramatically altering the LLM architecture. While providers may choose to alter the API to hide this information, the relevant API features have valuable and safe use cases for the LLM clients, who may rely on access to features like logit bias. Though our findings be viewed as a warning to LLM providers to carefully consider the consequences of their LLM architectures and APIs, we prefer to view our methods as a potential feature that LLM providers may choose to keep in order to better maintain trust with their customers by allowing outside observers to audit their model.

### 15 Mar 2024 v2

@@ Abstract
The commercialization of large language models (LLMs) has led to the common practice of restricting access to proprietary models via an limited API. In this work we show that, with a conservative assumption about the model architecture, it is possible to learn a surprisingly large amount of non-public information about an API-protected LLM from a small number of API queries (e.g., costing under \$1,000 for OpenAI’s gpt-3.5-turbo). Our findings are centered on one key observation: most modern LLMs suffer from a softmax bottleneck, which restricts the model outputs to a linear subspace of the full output space. We exploit this fact to unlock several capabilities: efficiently discovering the LLM’s hidden size, obtaining cheap full-vocabulary outputs, detecting and disambiguating different model updates, identifying the source LLM given a single full LLM output, and even estimating the output layer parameters. Our empirical investigations show the effectiveness of our methods, which allow us to estimate the embedding size of OpenAI’s gpt-3.5-turbo to be about 4,096. Lastly, we discuss ways that LLM providers can guard against these attacks, as well as how these capabilities can be viewed as a feature (rather than a bug) by allowing for greater transparency and accountability.
@@ 1 Introduction
Considering several proposals to mitigate this vulnerability, we find no obvious fix to prevent obtaining LLM images without dramatically altering the LLM architecture. While providers may choose to alter the API to hide this information, the relevant API features have valuable and safe use cases for the LLM clients, who may rely on access to features like logit bias. Though our findings be viewed as a warning to LLM providers to carefully consider the consequences of their LLM architectures and APIs, we prefer to view our findings as a potential feature that LLM providers may choose to keep in order to better maintain trust with their customers by allowing outside observers to audit their model.
@@ 10 Conclusion
11 Acknowledgements
We thank Jack Morris and Vitaly Shmatikov for their insights and feedback on this work.